OpenStudy (anonymous):

Find all solutions to sin x = 1/2√3 tan x in the interval [ 0 , 2pi ). How should I start to create like trig functions ?

OpenStudy (matheducatormcg):

substitute \[\tan(x)=\frac{\sin(x)}{\cos(x)}\]

OpenStudy (matheducatormcg):

let me know if you need another step

OpenStudy (anonymous):

Do I move the sin x to the right side by subtracting?

OpenStudy (matheducatormcg):

yes. then try to factor and use the ZPP(zero product property)

OpenStudy (anonymous):

do you mean this \[ \sin x =\frac{1}{2\sqrt{3}} \tan x\] or \[\sin x =\frac{1}{2\sqrt{3} \tan x}\]

OpenStudy (matheducatormcg):

good question. not sure. my answer has been assuming the first version you gave amir.

OpenStudy (anonymous):

It's the first version. @matheducatorMcG , is the cosine supposed to be 1/ cos x when you factored it & its on the demoninator? or it doesnt matter ?

OpenStudy (matheducatormcg):

this is what i meant \[0=\sin(x)(\frac{1}{2\sqrt{3}\cos(x)}-1)\]

OpenStudy (matheducatormcg):

that second factor seems suspect. \[\frac{1}{2\sqrt{3}\cos(x)}-1=0\] \[\frac{1}{2\sqrt{3}\cos(x)}=1\] \[1=2\sqrt{3}\cos(x)\] \[\frac{1}{2\sqrt{3}}=cos(x)\] I don't know that one off the top of my head. i would look at the graph of cos(x) and find intersections with \[y=\frac{1}{2\sqrt{3}}\]

OpenStudy (matheducatormcg):

2 intersections at \[x \approx1.2779536\] and \[x \approx 5.0052318\]

OpenStudy (matheducatormcg):

the other solutions for the sin(x)=0 were at 0, pi, and 2pi